A glass optical fiber is used to transport a light ray across a long distance. The fiber has an index of refraction of 1.550 and is submerged in glycerin, which has an index of refraction of 1.473. What is the critical angle (in degrees) for the light ray to remain inside the fiber?

°
Answer #1

Let n2 = 1.550

n1 = 1.473  

If Ic is the critical angle of incidence for a light to be inside the fibre, then

From Snell's law

Sin Ic  / sin 900 = n1/ n2

or, Sin Ic = 1.473/1.550

or, Ic = Sin-1( 1.473/1.550)

SO,  Ic = 71.860